# Weighted Pseudorandom Generators for Read-Once Branching Programs via Weighted Pseudorandom Reductions

### v2

@@ 5.2 The algorithm
There exists an algorithm that takes as input a regular ROBP and a parameter $\varepsilon>0$, and outputs an approximation of its acceptance probability with error $\varepsilon$. Furthermore, if the regular ROBP has length $n$, width $w$ and alphabet size $2^{s}$, then the algorithm runs in space $O(s+\log n(\log\log n+\sqrt{\log(w/\varepsilon)})+\log(w/\varepsilon))$.
Proof.
To define the algorithm, we need to prepare three things in advance: the two-way labeling, the error reduction, and the rotation algorithm.
First, the input of the algorithm is a regular ROBP $f$ without two-way labeling. However, we can easily convert it to a regular ROBP $G$ with two-way labeling by assigning an arbitrary label to each incoming edge. (i.e., $\ell_{in}(e)=i$ if $e$ is the $i$-th incoming edge of the vertex.). This can be done in space $O(s+\log(nw))$.
Second, we need to predetermine the sets of indices and signs used in a recursive error reduction procedure. We use parameters similar to those in Section 4. We set $n_{0}=n,n_{1}=O(\log n\sqrt{\log(w/\varepsilon)})$. For $i=2,3,\ldots$, let $n_{i}=\left\lceil\log^{3}n_{i-1}\right\rceil$. We take $l$ to be the first integer such that $n_{l}\leq 2^{15}$. Notice that $l=o(\log\log n)$.
Now for each $i\in[l]$, we prepare a group of parameters for invoking Theorem 5.13. We use $n_{i},n_{i+1}$ as the $n,r$ respectively in Theorem 5.13. When we apply it to $n_{0},n_{1}$, we set $\alpha_{0}=\varepsilon/2w$ as the target error and $k_{0}=\sqrt{\log(1/\alpha_{0})}$, $\tau_{0}=\min\{\frac{\varepsilon^{2/(k_{0}+1)}}{16\log^{2}n_{0}},\frac{1}{64\log^{2}n_{0}}\}$ as the original error. Theorem 5.13 gives a number $W_{0}\in\mathbb{N}$,a set of indices $\{n^{(0)}_{j,t}\}_{j\in[W_{0}],t\in[n_{1}]}$ and a set of signs $\{\sigma_{j}\}_{j\in[W_{0}]}$. After that, we set $\varepsilon^{\prime}=\varepsilon^{2}/(2wW_{0})$. Then for every $i\in[l-1]$, we set $\alpha_{i}=\varepsilon^{\prime}$ and $k_{i}=\log^{2}n_{i}$, $\tau_{i}=\min\{\frac{(\varepsilon^{\prime})^{2/(k_{i}+1)}}{16\log^{2}n_{i}},\frac{1}{64\log^{2}n_{i}}\}$. Theorem 5.13 gives a number $W_{i}\in\mathbb{N}$, a set of indices $\{n^{(i)}_{j,t}\}_{j\in[W_{i}],t\in[n_{i+1}]}$ and a set of signs $\{\sigma_{j}\}_{j\in[W_{i}]}$.
We stress that one can assume each $\{n^{(i)}_{j,t}\}_{j\in[W_{i}],t\in[n_{i+1}]}$ and $\{\sigma_{j}\}_{j\in[W_{i}]}$ satisfies the following properties:
For each $i=0,1,2,\ldots,l-1$, for any doubly stochastic matrices $A_{1},\ldots,A_{n_{i}}$ and $B_{j,q}$ such that $B_{p,q}\overset{sv}{\approx}_{\tau_{i}}A_{p+1}\ldots A_{q}$ and $B_{p-1,p}=A_{p}$ for all $p\in[n_{i}]$. Then let
$A^{\prime}=\sum_{j\in[W_{i}]}\sigma_{j}\cdot B_{0,n^{(i)}_{j,1}}B_{n^{(i)}_{j,1},n^{(i)}_{j,2}}\ldots B_{n^{(i)}_{j,n_{i+1}-1},n^{(i)}_{j,n_{i+1}}}.$
It holds that $A^{\prime}$ approximates $A_{1}\ldots A_{n_{i}}$ with entry-wise error at most $\alpha_{i}$, by Theorem 5.13.
The third part of the preparation is an implementation of the rotation algorithm Theorem 5.11. We use the algorithm for every level $i$, to get those approximation matrices with sv-error $\tau_{i}$.
For any $i=0,1,2,\ldots,l-1$, to use Lemma 5.10 for the correctness of Theorem 5.11, we need to prepare a group of parameters $n_{i},\lambda_{i},d_{i},c_{i}$. Let $\lambda_{i}$ be $\tau_{i}/(11\log n_{i})$. For each $t\in[\log n_{i}]$, we use $H^{(i)}_{t}$ as the $t$-th $\lambda_{i}$-spectral expander used in the lemma, where every $H^{(i)}_{t}$ can be achieved by Lemma 5.12. The parameter $c_{i}$ is set to be the degree of $H^{(i)}_{t}$, which is the same for all $t$. We let $d_{0}=2^{s}$. Let $d_{i+1}=d_{i}+\left\lceil\log n_{i}\right\rceil\cdot\left\lceil\log c_{i}\right\rceil$ for all $i=0,1,\ldots,l-1$. One can see later that the parameter $d_{i}$ is the degree of the branching program after the first $i$ levels of reduction.
Now we have prepared $l,\{n_{i}\}_{i=0}^{l},\{W_{i}\}_{i=0}^{l},\{n^{(i)}_{j,t}\}_{i\in[l],j\in[W_{i}],t\in[n_{i+1}]},\{\sigma_{j}\}_{j\in[n_{i}]},\{d_{i}\}_{i=0}^{l},\{c_{i}\}_{i=1}^{l}$ and graphs $\{H^{(i)}_{t}\}_{i\in[l],t\in[\log n_{i}]}$. Notice that all these parameters can be computed in logspace, so we will directly use them in the algorithm description. Then we construct our algorithm that approximates the acceptance probability of the regular ROBP. In general the algorithm adapts the WPRG of the previous section and runs the ROBP on the output of the WPRG but with the following two adjustments:
  • every time when the WPRG needs to conduct a one-step walk on an expander, then the algorithm instead conducts a corresponding rotation on that expander.
  • every time when running one step of the ROBP using an output symbol of the WPRG, the algorithm conducts a corresponding rotation on the ROBP.
For completeness of the description, the whole algorithm is given as Algorithm 1.
In Algorithm 1, for every $j\in[n]$, $G_{j\leftarrow j-1}$ is the bipartite graph $(V_{j-1},V_{j},E_{j})$ of input ROBP $f$, as described in Definition 5.9. $\mathsf{Rot}_{H^{(i)}_{p}}$ and $\mathsf{Rot}_{G_{j\leftarrow j-1}}$ are defined by Definition 5.4. We use the $x^{(i)}$ to denote a temporary variable that stores a label in the following way. $x^{(l)}\in[d_{l}]=[d_{0}]\times[c_{0}]^{\log n_{0}}\times[c_{1}]^{\log n_{1}}\times\cdots\times[c_{l-1}]^{\log n_{l-1}}$. For every $i\in\{0,1,...,l-1\}$, $x^{(i)}$ denotes the prefix of $x^{(l)}$ in $[d_{i}]=[d_{0}]\times[c_{0}]^{\log n_{0}}\times\cdots\times[c_{i-1}]^{\log n_{i-1}}$, and $e^{(i)}_{t}$ is in $[c_{i}]$ for every $t\in[\log n_{i}]$ such that $x^{(i)}=(x^{(i-1)},e^{(i-1)}_{1},...,e^{(i-1)}_{\log n_{i-1}})$.
Input: A regular ROBP $f$, a starting vertex $v_{0}$, an accepting vertex $v_{end}$, an error parameter $\varepsilon>0$.
Output: An approximation of the probability that $f$ reaches $v_{end}$ from $v_{0}$.
Set a counter $ans\leftarrow 0$.
foreach $j_{0}\in[W_{0}],j_{1}\in[W_{1}],\ldots,j_{l-1}\in[W_{l-1}]$(Enumerate all possible summands in the error reduction polynomial) do
       foreach $(x^{(l)}_{1},x^{(l)}_{2},\ldots,x^{(l)}_{n_{l}})$ (Enumerate seeds), where each $x^{(l)}_{t}\in[d_{l}]=[d_{0}]\times[c_{0}]^{\log n_{0}}\times[c_{1}]^{\log n_{1}}\times\cdots\times[c_{l-1}]^{\log n_{l-1}}$ do
             Let $x^{(l)}$ be a temporary variable and $x^{(l)}\leftarrow x^{(l)}_{1}$.
            Initiate position pointers for each level $t_{0}\leftarrow 0,t_{1}\leftarrow 0,\ldots,t_{l}\leftarrow 0$.
            Let $v\leftarrow v_{0}$ be a temporary variable recording the current vertex.
            while $t_{l}<n_{l}$ do
                   $t_{0}\leftarrow t_{0}+1$.
                  $(v,x^{(0)})\leftarrow\mathsf{Rot}_{G_{t_{0}\leftarrow t_{0}-1}}(v,x^{(0)})$.
                  Update pointers: For $i=1$ to $l$, if $t_{i-1}=n^{(i-1)}_{j_{i-1},t_{i}+1}$, we let $t_{i}\leftarrow t_{i}+1$. Let $i$ be the largest $i$ such that $t_{i}$ is updated.
                  if $i<l$ then
                        Let $p\in[\log n_{i}]$ be the smallest positive integer such that $t_{i}-n^{(i)}_{j_{i},t_{i+1}}$ is not a multiple of $2^{p}$.
                        Update $(x^{(i)},e^{(i)}_{1},e^{(i)}_{2},\ldots,e^{(i)}_{p-1})\leftarrow\mathsf{Rot}_{H^{(i)}_{p}}((x^{(i)},e^{(i)}_{1},e^{(i)}_{2},\ldots,e^{(i)}_{p-1}),e^{(i)}_{p})$.
                  else if $i=l$ then
                         $x^{(l)}\leftarrow x^{(l)}_{t_{l}+1}$
            if $v=v_{end}$ then
                   Add $\sigma_{j_{0}}\cdot\sigma_{j_{1}}\cdots\sigma_{j_{l-1}}$ to $ans$.
Normalize $ans$ by $ans\leftarrow ans/d_{l}$, return $ans$.
Algorithm 1 Approximate the acceptance probability of a regular ROBP
The following figure gives an example of how our algorithm 1 computes the desired rotation maps.
$\mathsf{Rot}_{\widetilde{G}^{(2)}_{0\to 1}}$$\mathsf{Rot}_{\widetilde{G}^{(2)}_{1\to 2}}$$\mathsf{Rot}_{\widetilde{G}^{(1)}_{0\to 1}}$$\mathsf{Rot}_{\widetilde{G}^{(1)}_{1\to 2}}$$\mathsf{Rot}_{\widetilde{G}^{(1)}_{2\to 3}}$$\mathsf{Rot}_{\widetilde{G}^{(1)}_{3\to 4}}$$\mathsf{Rot}_{\widetilde{G}^{(1)}_{4\to 5}}$$\mathsf{Rot}_{\widetilde{G}^{(0)}_{0\to 1}}$$\mathsf{Rot}_{\widetilde{G}^{(0)}_{1\to 2}}$$\mathsf{Rot}_{\widetilde{G}^{(0)}_{2\to 3}}$$\mathsf{Rot}_{\widetilde{G}^{(0)}_{3\to 4}}$$\mathsf{Rot}_{\widetilde{G}^{(0)}_{4\to 5}}$$\mathsf{Rot}_{\widetilde{G}^{(0)}_{5\to 6}}$$\mathsf{Rot}_{\widetilde{G}^{(0)}_{6\to 7}}$$\mathsf{Rot}_{\widetilde{G}^{(0)}_{7\to 8}}$$\mathsf{Rot}_{\widetilde{G}^{(0)}_{8\to 9}}$$\mathsf{Rot}_{\widetilde{G}^{(0)}_{9\to 10}}$$\mathsf{Rot}_{\widetilde{H}^{(0)}_{1}}$$\mathsf{Rot}_{\widetilde{H}^{(0)}_{1}}$$\mathsf{Rot}_{\widetilde{H}^{(0)}_{1}}$$\mathsf{Rot}_{\widetilde{H}^{(0)}_{1}}$$\mathsf{Rot}_{\widetilde{H}^{(0)}_{1}}$$\mathsf{Rot}_{\widetilde{H}^{(1)}_{1}}$$\mathsf{Rot}_{\widetilde{H}^{(1)}_{1}}$$\mathsf{Rot}_{\widetilde{H}^{(1)}_{2}}$se the next seed $x^{(2)}_{2}$se seed $x^{(2)}_{1}$
Figure 3: The figure shows our recursion process to compute $\mathsf{Rot}_{\widetilde{G}^{(l)}_{1\to 2}}\left(\mathsf{Rot}_{\widetilde{G}^{(l)}_{0\to 1}}\left(v_{0},x^{(l)}_{1}\right),x^{(l)}_{2}\right)$ when $l=2,n_{l}=2$. The indices $j_{0},j_{1},j_{2}$ are omitted.
Next we prove the correctness of the algorithm.
Let $\widetilde{G}^{(0)}_{t-1\to t}$ denote the bipartite graph with two-way labeling of the input from the $(t-1)$-th layer to the $t$-th layer. We recursively define $\widetilde{G}^{(i)}_{j_{0},j_{1},...,j_{i-1},t-1\to t}$ as the graph obtained by applying Theorem 5.11 to $\widetilde{G}^{(i-1)}_{j_{0},j_{1},...,j_{i-2},a\to a+1},\ldots,\widetilde{G}^{(i-1)}_{j_{0},j_{1},...,j_{i-2},b-1\to b}$ where $b=n^{(i)}_{j_{i-1},t},a=n^{(i)}_{j_{i-1},t-1}$ with the aforementioned parameters $n_{i},\lambda_{i},d_{i},c_{i}$. Let $\widetilde{\mathbf{W}}^{(i)}_{j_{0},j_{1},...,j_{i-1},t\leftarrow t-1}$ be the transition matrix of $\widetilde{G}^{(i)}_{j_{0},j_{1},...,j_{i-1},t-1\to t}$.
The output of Algorithm 1 is a weighted sum of $\widetilde{\mathbf{W}}^{(l)}_{j_{0},j_{1},...,j_{i-1},n_{l}\leftarrow n_{l}-1}\cdots\widetilde{\mathbf{W}}^{(l)}_{j_{0},j_{1},...,j_{i-1},1\leftarrow 0}$, as shown in the following equation:
$ans$
$=$ $\sum_{\begin{subarray}{c}j_{0}\in[W_{0}],j_{1}\in[W_{1}],\\ \ldots,j_{l-1}\in[W_{l-1}]\end{subarray}}\left(\prod_{i=0}^{l-1}\sigma_{j_{i}}\right)\cdot\mathbb{E}_{x^{(l)}_{1},x^{(l)}_{2},\ldots,x^{(l)}_{n_{l}}\in[d_{l}]}$
$\left[\mathbf{1}\left[\mathsf{Rot}_{\widetilde{G}^{(l)}_{\begin{subarray}{c}j_{0},j_{1},\ldots,j_{l-1}\\ ,n_{l}-1\to n_{l}\end{subarray}}}\left(\cdots\mathsf{Rot}_{\widetilde{G}^{(l)}_{\begin{subarray}{c}j_{0},j_{1},\cdots,j_{l-1}\\ ,1\to 2\end{subarray}}}\left(\mathsf{Rot}_{\widetilde{G}^{(l)}_{\begin{subarray}{c}j_{0},j_{1},\cdots,j_{l-1}\\ ,0\to 1\end{subarray}}}\left(v_{0},x^{(l)}_{1}\right),x^{(l)}_{2}\right)\ldots,x^{(l)}_{n_{l}}\right)=v_{end}\right]\right]$
$=$ $\sum_{\begin{subarray}{c}j_{0}\in[W_{0}],j_{1}\in[W_{1}],\\ \ldots,j_{l-1}\in[W_{l-1}]\end{subarray}}\left(\prod_{i=0}^{l-1}\sigma_{j_{i}}\right)\cdot\left[\widetilde{\mathbf{W}}^{(l)}_{j_{0},j_{1},...,j_{i-1},n_{l}\leftarrow n_{l}-1}\cdots\widetilde{\mathbf{W}}^{(l)}_{j_{0},j_{1},...,j_{i-1},1\leftarrow 0}\right]_{v_{0},v_{end}}.$
By Lemma 5.10, for any $i\in[l],j_{0},j_{1},...,j_{i-1}\in[W_{0}],[W_{1}],...,[W_{i-1}]$, letting $b=n^{(i)}_{j_{i-1},t},a=n^{(i)}_{j_{i-1},t-1}$, we have the following approximation.
$\widetilde{\mathbf{W}}^{(i)}_{j_{0},j_{1},...,j_{i-1},t\leftarrow t-1}\overset{sv}{\approx}_{\tau_{i-1}}\widetilde{\mathbf{W}}^{(i-1)}_{j_{0},j_{1},...,j_{i-2},b\leftarrow b-1}\widetilde{\mathbf{W}}^{(i-1)}_{j_{0},j_{1},...,j_{i-2},b-1\leftarrow b-2}\ldots\widetilde{\mathbf{W}}^{(i-1)}_{j_{0},j_{1},...,j_{i-2},a\leftarrow a+1}.$
These matrices satisfy the condition of Theorem 5.13. By the construction of the $\{n^{(i)}_{j,t}\}$ and $\{\sigma_{j}\}$, we have that for each $i\in{0,1,\ldots,l-1}$ and each $j_{0},j_{1},\ldots,j_{i-1}$,
$\left|\left[\widetilde{\mathbf{W}}^{(i)}_{\begin{subarray}{c}j_{0},...,j_{i-1}\\ n_{i}\leftarrow n_{i}-1\end{subarray}}\cdots\widetilde{\mathbf{W}}^{(i)}_{\begin{subarray}{c}j_{0},...,j_{i-1}\\ 1\leftarrow 0\end{subarray}}\right]_{v_{0},v_{end}}-\sum_{j_{i}\in[W_{i}]}\left[\sigma_{j_{i}}\cdot\widetilde{\mathbf{W}}^{(i+1)}_{\begin{subarray}{c}j_{0},...,j_{i}\\ n_{i+1}\leftarrow n_{i+1}-1\end{subarray}}\cdots\widetilde{\mathbf{W}}^{(i+1)}_{\begin{subarray}{c}j_{0},...,j_{i}\\ 1\leftarrow 0\end{subarray}}\right]_{v_{0},v_{end}}\right|\leq\alpha_{i}.$
By summing over all the errors, we have the following:
$\left|ans-\left[\widetilde{\mathbf{W}}^{(0)}_{n_{0}\leftarrow n_{0}-1}\cdots\widetilde{\mathbf{W}}^{(0)}_{1\leftarrow 0}\right]_{v_{0},v_{end}}\right|$
$=$ $\left|\sum_{i=0}^{l-1}\sum_{\begin{subarray}{c}j_{0}\in[W_{0}],j_{1}\in[W_{1}],\\ \ldots,j_{i-1}\in[W_{i-1}]\end{subarray}}\left(\prod_{k=0}^{i-1}\sigma_{j_{k}}\right)\right.$
[…]

### v3

@@ 5.2 The derandomization for regular ROBPs with large alphabets
There exists an algorithm that takes as input a regular ROBP and a parameter $\varepsilon>0$, and outputs an approximation of the acceptance probability with additive error $\varepsilon$. Furthermore, if the regular ROBP has length $n$, width $w$ and alphabet bit length $s$, then the algorithm uses $O(s+\log n(\log\log n+\sqrt{\log(w/\varepsilon)})+\log(w/\varepsilon))$ bits of space.
Let $f\in\mathcal{R}^{tw}_{n,w,s}$ be a regular ROBP with two-way labeling and let $\mathbf{M}_{l\ldots r}$ denote its transition matrix from layer $l$ to layer $r$. For every $0\leq l<r\leq n$, we invoke Lemma 5.12, which gives a regular bigraph $\widetilde{G}_{l\ldots r}$ with two-way labeling such that its transition matrix $\widetilde{\mathbf{M}}_{l\ldots r}$ is a $\tau$-sv-approximation of $\mathbf{M}_{l\ldots r}$. Then by Theorem 5.16, for any $k\in\mathbb{N}$, we have $K=O((2n)^{k})$ and a partition $0=m_{i,0}<m_{i,1}<\cdots<m_{i,l}=n_{0}$ for every $i\in[K]$ and $\sigma_{i}$ such that
$\sum_{i\in[K]}\sigma_{i}\widetilde{\mathbf{M}}_{m_{i,0}\ldots m_{i,1}}\widetilde{\mathbf{M}}_{m_{i,1}\ldots m_{i,2}}\cdots\widetilde{\mathbf{M}}_{m_{i,l-1}\ldots m_{i,l}}\text{ approximates }\mathbf{M}_{0\ldots n_{0}}\text{ with entry-wise error }\tau^{k},$
Given the summation, we can view it as a sum of regular ROBPs with two-way labelings. For every $i\in[K]$, we define $f_{i}\in\mathcal{R}^{tw}_{n_{1},w,s_{1}}$ as the concatenation of the bipartite graphs
$\widetilde{G}_{m_{i,0}\ldots m_{i,1}},\widetilde{G}_{m_{i,1}\ldots m_{i,2}},\ldots,\widetilde{G}_{m_{i,l-1}\ldots m_{i,l}},$
which still has a two-way labeling. We set the starting vertex and accepting vertices of $f_{i}$ to be the same as those of $f$. Then we obtain
$\left|\mathbb{E}_{x}f(x)-\sum_{i\in[K]}\sigma_{i}\mathbb{E}_{y}f_{i}(y)\right|\leq\tau^{k}w.$
The same procedure can be applied to each regular ROBP $f_{i}$, which gives shorter ROBPs $f_{i_{1},i_{2}}$ for every $i_{2}\in[K_{2}]$. We repeate the procedure until we attain regular ROBPs $f_{i_{1},\ldots,i_{\ell}}$ with constant length, which can be fooled by true randomness. We formalize the above into the following detailed proof.
Proof of Theorem 5.17.
Let $f\in\mathcal{R}^{tw}_{n,w,s}$ be the regular ROBP to derandomize. Let $\ell\in\mathbb{N}$ be the number of iterations we need to perform and let $\tau_{i},\varepsilon_{i}$ be the error parameters of the $i$-th iteration for every $i\in[\ell]$. The exact values of ,$\ell,\{\tau_{i},\varepsilon_{i}\}_{i\in[\ell]}$ will be specified later. For each iteration $i\in[\ell]$, define the error reduction exponent $k_{i}=\lceil\log_{1/\tau_{i}}(2/\varepsilon_{i})\rceil$, and we setup $n_{i},K_{i}$ iteratively. Let $n_{0}=n$. Consider every $i\in[\ell]$. Let $n_{i}=t,K_{i}=K$, where $t=O(k_{i}\cdot\log n_{i-1})$ and $K=O((2n_{i-1})^{k_{i}})$ are the parameters whose existence is guaranteed by Theorem 5.16, when initiating Theorem 5.16 with $n=n_{i-1},k=k_{i},\tau=\tau_{i}$.
We iteratively define $f_{i_{1},i_{2},\ldots,i_{p}}\in\mathcal{R}^{tw}_{n_{p},w,s_{p}}$ for every $i_{1}\in[K_{1}],i_{2}\in[K_{2}],\ldots,i_{p}\in[K_{p}]$ and $p\in\{0,1,\ldots,\ell\}$ as follows:
  1. 1.
    When $p=0$, $f_{i_{1},i_{2},\ldots,i_{p}}$ is defined to be $f$.
  2. 2.
    For every $p\in[\ell]$, assume we have defined $f_{i_{1},i_{2},\ldots,i_{p-1}}$.We denote $\mathbf{M}^{(i_{1},i_{2},\ldots,i_{p-1})}_{l\ldots r}$ to be the stochastic matrix of $f_{i_{1},i_{2},\ldots,i_{p-1}}$ from layer $l$ to layer $r$. For every $0\leq l<r\leq n_{p-1}$, let $\widetilde{G}^{(i_{1},i_{2},\ldots,i_{p-1})}_{l\ldots r}$ be a regular bigraph with a two-way labeling obtained by approximating $\mathbf{M}^{(i_{1},i_{2},\ldots,i_{p-1})}_{l\ldots r}$ by the algorithm of Lemma 5.12 with error parameter $\tau_{p}$ 777we always consider it as a length $2^{\left\lceil\log n_{p-1}\right\rceil}$ ROBP regardless of $r-l$. Let $\widetilde{\mathbf{M}}^{(i_{1},i_{2},\ldots,i_{p-1})}_{l\ldots r}$ be the stochastic matrix of $\widetilde{G}^{(i_{1},i_{2},\ldots,i_{p-1})}_{l\ldots r}$. By Theorem 5.16, for every $i_{p}\in[K_{p}]$ we have a partition $0=m^{(p)}_{i_{p},0}<m^{(p)}_{i_{p},1}<\cdots<m^{(p)}_{i_{p},n_{p}}=n_{p-1}$ and weights $\sigma^{(p)}_{i_{p}}\in\{-1,0,1\}$ such that
    $\sum_{i_{p}\in[K_{p}]}\sigma^{(p)}_{i_{p}}\widetilde{\mathbf{M}}^{(i_{1},i_{2},\ldots,i_{p-1})}_{m^{(p)}_{i_{p},0}\ldots m^{(p)}_{i_{p},1}}\widetilde{\mathbf{M}}^{(i_{1},i_{2},\ldots,i_{p-1})}_{m^{(p)}_{i_{p},1}\ldots m^{(p)}_{i_{p},2}}\cdots\widetilde{\mathbf{M}}^{(i_{1},i_{2},\ldots,i_{p-1})}_{m^{(p)}_{i_{p},n_{p}-1}\ldots m^{(p)}_{i_{p},n_{p}}}\text{ approximates }$ (1)
    $\mathbf{M}^{(i_{1},i_{2},\ldots,i_{p-1})}_{0\ldots n_{p-1}}\text{with entry-wise error }\tau_{p}^{k_{p}}.$
    We define $f_{i_{1},i_{2},\ldots,i_{p}}$ to be the concatenation of $\widetilde{G}^{(i_{1},i_{2},\ldots,i_{p-1})}_{m^{(p)}_{i_{p},0}\ldots m^{(p)}_{i_{p},1}},\widetilde{G}^{(i_{1},i_{2},\ldots,i_{p-1})}_{m^{(p)}_{i_{p},1}\ldots m^{(p)}_{i_{p},2}},\ldots,\widetilde{G}^{(i_{1},i_{2},\ldots,i_{p-1})}_{m^{(p)}_{i_{p},n_{p}-1}\ldots m^{(p)}_{i_{p},n_{p}}}$, which also has a two-way labeling.
Having defined the regular ROBPs, we use $|\Sigma_{p}|=2^{s_{p}}$ to denote the alphabet size of $f_{i_{1},i_{2},\ldots,i_{p}}$. Note that for every $i_{1}\in[K_{1}],i_{2}\in[K_{2}],\ldots,i_{p}\in[K_{p}]$ and $p\in\{0,1,\ldots,\ell\}$, it holds that $f_{i_{1},i_{2},\ldots,i_{p}}$ always has the same alphabet size, since all such regular ROBPs are constructed from Theorem 5.13 with the same set of parameters.
Our derandomization algorithm computes
$\sum_{i_{1}\in[K_{1}],i_{2}\in[K_{2}],\ldots,i_{\ell}\in[K_{\ell}]}\sigma^{(1)}_{i_{1}}\sigma^{(2)}_{i_{2}}\cdots\sigma^{(\ell)}_{i_{\ell}}\cdot\mathbb{E}_{x\in(\{0,1\}^{s_{\ell}})^{n_{\ell}}}f_{i_{1},i_{2},\ldots,i_{\ell}}(x).$
The algorithm operates as follows, assuming we can access the ROBP $f_{i_{1},i_{2},\ldots,i_{\ell}}$ with two-way labeling:
  1. 1.
    Enumerate all tuples $(i_{1},i_{2},\ldots,i_{\ell})\in[K_{1}]\times[K_{2}]\times\cdots\times[K_{\ell}]$ and every string $(x_{1},x_{2},\ldots,x_{n_{\ell}})\in(\Sigma_{\ell})^{n_{\ell}}$. For each tuple $(i_{1},i_{2},\ldots,i_{\ell},x_{1},x_{2},\ldots,x_{n_{\ell}})$, we compute the following:
    1. (a)
      Let $u\in[w]$ be the starting vertex of $f_{i_{1},i_{2},\ldots,i_{\ell}}$.
    2. (b)
      For every $t\in[n_{\ell}]$, we compute the rotation map $u\leftarrow\mathsf{Rot}_{f_{i_{1},i_{2},\ldots,i_{\ell}}}(u,x_{t})$. ($\mathsf{Rot}$ outputs a new pair $(u,x_{t})$, but we discard $x_{t}$.)
    3. (c)
      If $u$ is an accepting vertex of $f_{i_{1},i_{2},\ldots,i_{\ell}}$, then add $\sigma^{(1)}_{i_{1}}\sigma^{(2)}_{i_{2}}\cdots\sigma^{(\ell)}_{i_{\ell}}/(2^{n_{l}\cdot s_{l}})$ to the result; otherwise, add 00 to the result.
  2. 2.
    Return the result.
Since we do not have a direct access to $f_{i_{1},i_{2},\ldots,i_{\ell}}$, we need to compute it with a recursive program. For every $p\in[\ell]$ all tuples $(i_{1},i_{2},\ldots,i_{p})\in[K_{1}]\times[K_{2}]\times\cdots\times[K_{p}]$ we compute $(u^{\prime},x^{\prime})\leftarrow\mathsf{Rot}_{f_{i_{1},i_{2},\ldots,i_{p}}}(u,x)$ with the following recursive program, under the assumption that we have already defined the program to compute $\mathsf{Rot}_{f_{i_{1},i_{2},\ldots,i_{p-1}}}(\cdot,\cdot)$:
  1. 1.
    If $p=0$, then we compute $\mathsf{Rot}_{f}(u,x)$ directly with the original ROBP $f$.
  2. 2.
    Otherwise, assume $u=(a,b)$ is the $a$-th vertex in the $b$-th layer of the ROBP $f_{i_{1},i_{2},\ldots,i_{p}}$. Then we let $b^{\prime}\leftarrow b+1$ and compute $(a^{\prime},x^{\prime})\leftarrow\mathsf{Rot}_{\widetilde{G}^{(i_{1},i_{2},\ldots,i_{p-1})}_{m^{(p)}_{i_{p},b}\ldots m^{(p)}_{i_{p},b+1}}}(a,x)$ with $\operatorname{DerandWalk}$ from Definition 5.15:
    1. (a)
      Let $l\leftarrow m^{(p)}_{i_{p},b},r\leftarrow m^{(p)}_{i_{p},b+1}$.
    2. (b)
      Let $(a^{\prime},x^{\prime})\leftarrow\operatorname{DerandWalk}(\tau_{p},f_{i_{1},i_{2},\ldots,i_{p-1}},l,r,a,x)$, during which the algorithm calls $\mathsf{Rot}_{f_{i_{1},i_{2},\ldots,i_{p-1}}}$.
    3. (c)
      Let $u^{\prime}\leftarrow(a^{\prime},b^{\prime})$.
    4. (d)
      Output $(u^{\prime},x^{\prime})$.
We now examine both the error bound and the complexity of the algorithm. First we bound the error of the algorithm. By the construction of $f_{i_{1},i_{2},\ldots,i_{p}}$, we have
$\left|\mathbb{E}_{x}f_{i_{1},i_{2},\ldots,i_{p-1}}(x)-\sum_{i_{p}\in[K_{p}]}\sigma^{(p)}_{i_{p}}\mathbb{E}_{y}f_{i_{1},i_{2},\ldots,i_{p}}(y)\right|$ $\leq\tau_{p}^{k_{p}}w\leq\varepsilon_{p}w$
for every $i_{1}\in[K_{1}],i_{2}\in[K_{2}],\ldots,i_{p-1}\in[K_{p-1}]$. Using triangular inequalities, we have the following bound:
$\left|\mathbb{E}_{x}f(x)-\sum_{i_{1}\in[K_{1}],i_{2}\in[K_{2}],\ldots,i_{\ell}\in[K_{\ell}]}\sigma^{(1)}_{i_{1}}\sigma^{(2)}_{i_{2}}\cdots\sigma^{(\ell)}_{i_{\ell}}\cdot\mathbb{E}_{y}f_{i_{1},i_{2},\ldots,i_{\ell}}(y)\right|$
$\leq\sum_{p=0}^{\ell-1}\sum_{i_{1}\in[K_{1}],i_{2}\in[K_{2}],\ldots,i_{p}\in[K_{p}]}\left|\mathbb{E}_{x}f_{i_{1},i_{2},\ldots,i_{p}}(x)-\sum_{i_{p+1}\in[K_{p+1}]}\sigma^{(p+1)}_{i_{p+1}}\mathbb{E}_{y}f_{i_{1},i_{2},\ldots,i_{p},i_{p+1}}(y)\right|$
$\leq\varepsilon_{1}w+K_{1}\cdot\varepsilon_{2}w+K_{1}K_{2}\cdot\varepsilon_{3}w+\cdots+K_{1}K_{2}\cdots K_{\ell-1}\cdot\varepsilon_{\ell}w$
For the complexity of the algorithm, notice that the recursive algorithm computing $\mathsf{Rot}_{f_{i_{1},i_{2},\ldots,i_{p}}}$ requires $(s_{p}-s_{p-1})+\log n_{p}$ more bits than that of the algorithm computing $\mathsf{Rot}_{f_{i_{1},i_{2},\ldots,i_{p-1}}}$, since it only needs to store $b_{l},b_{r}$ and the current $e_{1},e_{2},\ldots,e_{\log\left\lceil n_{p-1}\right\rceil}$ during the computation of $((a,b),y)\leftarrow\mathsf{Rot}_{f_{i_{1},i_{2},\ldots,i_{p-1}}}((a,b),y)$. Given that the base case $p=0$ requires $O(\log n_{0}+s_{0}+\log w)$ bits of space, the algorithm to compute $\mathsf{Rot}_{f_{i_{1},i_{2},\ldots,i_{\ell}}}$ requires $O(\sum_{p=0}^{\ell}\log n_{p}+s_{\ell}+\log w)$ bits of space. The main algorithm to compute the expectation of $f_{i_{1},i_{2},\ldots,i_{\ell}}$ requires $O(s_{\ell}+\sum_{p=0}^{\ell}\log K_{p}+\log w)$ bits to store the current tuple $(i_{1},i_{2},\ldots,i_{\ell})$ and the current string $(x_{1},x_{2},\ldots,x_{n_{\ell}})$ and maintain the result. Therefore, the total space complexity of the algorithm is
$O\left(\sum_{p=0}^{\ell}\log n_{p}+s_{\ell}+\log w+\sum_{p=0}^{\ell}\log K_{p}\right).$
[…]